# fixedprepend, caption=FIXED, color=green

### v2

@@ 4 Experimental Results
We first evaluate ChemDFM-R with multiple baselines on several benchmarks specifically designed for assessing the chemical capabilities of LLMs (§˜4.1). Then, to highlight the importance and effectiveness of chemical reasoning, we analyze the rationale of our model in detail (§˜4.2) and demonstrate its capability in facilitating reliable human-AI collaboration (§˜4.3).
@@ 4.1 Benchmark Evaluation
We evaluate ChemDFM-R and the baseline models on two of the most popular and comprehensive benchmarks specifically designed for assessing the chemical capabilities of LLMs: SciKnowEval (Feng et al., 2024) and ChemEval (Huang et al., 2024). Given the large number of tasks included in SciKnowEval (19 tasks) and ChemEval (36 tasks), to facilitate fair and clear comparison, we categorized the tasks into three groups: text-centric, molecule-centric, and reaction-centric tasks.
@@ 4.1.1 Performance Comparison with Baselines
First, we demonstrate the effectiveness of our training pipeline by comparing the performances of ChemDFM-R with those of 1) Qwen2.5-14B-Instruct (Yang et al., 2024a), which is the general-domain instruction tuning model of Qwen2.5-14B, and 2) ChemDFM-I, which incorporates atomized chemical knowledge enhancement but precedes the stage of chemical rationale learning. The results are illustrated in Table 1.
Table 1: Benchmark results on SciKnowEval and ChemEval. “mol.” stands for “molecule” and “react.” stands for “reaction”. The best performance for each task is indicated using boldface. * We use PRS (Peng et al., 2025) to balance the different scales of the scores on different tasks in the ChemEval benchmark.
@@
Qwen2.5-14B-Instruct 0.77 0.36 0.72 0.61 0.78 0.24 0.47 0.57
ChemDFM-I 0.77 0.50 0.95 0.69 0.80 0.69 0.60 0.72
ChemDFM-R 0.77 0.52 0.95 0.70 0.80 0.88 0.66 0.78
As showcased in Table 1, ChemDFM-R consistently outperforms Qwen2.5-14B-Instruct on both SciKnowEval and ChemEval. This demonstrates that our specialization pipeline has successfully improved the model’s chemical capabilities. Specifically, the performances on text-centric tasks remain almost intact while those on molecule-centric and reaction-centric tasks increase significantly, leading to a large boost in the overall performance. This demonstrates that our method manages to improve the chemical capabilities of LLM while largely maintaining its abilities in understanding natural language.
When taking the ChemDFM-I model into consideration, we observe a clear performance improvement trend from Qwen2.5-14B-Instruct to ChemDFM-I and further to ChemDFM-R. This demonstrates that both the atomized chemical knowledge enhancement stage and the chemical rationale learning stage are crucial in achieving the final performance improvement. The atomized chemical knowledge enhancement stage enhances the model’s chemical knowledge, while the chemical rationale learning stage teaches the model how to conduct plausible reasoning with this knowledge. Beyond improved performance, the chemical reasoning capability of ChemDFM-R also enhances reliability and facilitates more effective human–AI collaboration, which will be discussed in subsequent sections.
Furthermore, Figure 3 illustrates the performance changes across individual tasks. The results clearly show that most tasks benefit from our training pipeline, especially the molecule-centric tasks and reaction-centric tasks. Moreover, the two training stages provide complementary gains across different tasks, enabling the final model to achieve superior results on a broader range of tasks. Notably, among the tasks where ChemDFM-R does not surpass Qwen2.5-14B-Instruct, a substantial proportion involves numerical prediction, such as Yield Extraction, Molecular Property Calculation, and Product Yield Prediction. In fact, all the molecule-centric and reaction-centric tasks where ChemDFM-R falls short of Qwen2.5-14B-Instruct are those involving numerical reasoning and prediction. This pattern suggests that the numerical calculation and prediction abilities of ChemDFM-R are relatively weak, highlighting a potential direction for further improvements.
@@ 4.1.2 Performance Comparison with Advanced LLMs
To further demonstrate the prowess of ChemDFM-R, we compare it with three sets of models: 1) existing chemical LLMs, including MolInst (Fang et al., 2024), ChemLLM (Zhang et al., 2024a), and ChemDFM (Zhao et al., 2025b); 2) advanced non reasoning LLMs in the general domain, including GPT-4o (Hurst et al., 2024) and Qwen3-14B (no think) (Yang et al., 2025); 3) advanced reasoning LLMs in the general domain, including DeepSeek-R1 (Guo et al., 2025), Qwen3-14B (think) (Yang et al., 2025), and o4-mini (OpenAI, 2025a). The experimental results are illustrated in Table 2. For detailed task categorization and the performance of individual tasks, please refer to Appendix E.
Table 2: Benchmark results on SciKnowEval and ChemEval. “mol.” stands for “molecule” and “react.” stands for “reaction”. The best performance for each task is indicated using boldface, while the second-best is indicated using underline. * We use PRS (Peng et al., 2025) to balance the different scales of the scores on different tasks in the ChemEval benchmark.
@@
MolInst 0.69 0.38 0.42 0.55 0.47 0.16 0.10 0.29
ChemLLM-20B-DPO 0.74 0.38 0.88 0.62 0.65 0.25 0.19 0.43
ChemDFM-13B-v1.0 0.70 0.43 0.85 0.62 0.62 0.43 0.36 0.50
ChemDFM-8B-v1.5 0.72 0.46 0.87 0.64 0.27 0.13 0.19 0.21
GPT-4o 0.76 0.45 0.45 0.61 0.83 0.36 0.56 0.65
Qwen3-14B (no think) 0.77 0.35 0.80 0.61 0.82 0.25 0.28 0.53
DeepSeek-R1 0.80 0.31 0.83 0.62 0.81 0.20 0.50 0.58
Qwen3-14B (think) 0.87 0.41 0.88 0.70 0.79 0.25 0.43 0.56
o4-mini 0.81 0.57 0.97 0.74 0.78 0.63 0.59 0.69
ChemDFM-R 0.77 0.52 0.95 0.70 0.80 0.88 0.66 0.78
As shown in Table 2, ChemDFM-R significantly outperforms both the general-domain LLMs and domain-specific LLMs of similar size, especially in the molecule-centric and reaction-centric tasks. Specifically, it even surpasses Qwen3-14B, the next-generation model in the same series as our base model Qwen2.5-14B. When compared to cutting-edge LLMs, ChemDFM-R achieves better performance than GPT-4o and DeepSeek-R1, while demonstrating competitive results relative to o4-mini. Considering the tiny size of our model, this result strongly demonstrates the prowess of ChemDFM-R and the effectiveness of our specialization process.
4.2 Rationale Analysis
To further validate the correctness and precision of ChemDFM-R’s reasoning capability, we conducted a thorough analysis of the rationales generated by ChemDFM-R. An example is illustrated in Figure 4, with additional examples provided in the Appendix F.
@@ 4.3 Reliable Human-AI Collaboration
Another important advantage of ChemDFM-R’s reasoning capability is that it allows humans to verify the correctness of answers, identify and correct errors in the reasoning process, and discover new insights or perspectives. This enables more practical, reliable, and flexible human-AI collaboration. In this section, we demonstrate the prowess of ChemDFM-R to facilitate reliable human-AI collaboration through examples. Specifically, an example is showcased in Figure 5, while more examples are illustrated in Appendix G.
@@
Figure 5: An example of reliable human-AI collaboration using ChemDFM-R. We draw inspiration from this conversation from Li et al. (2025). This example illustrates a process of research inspiration seeking with the help of ChemDFM-R.
@@
Specifically, to make our example more realistic, we design our conversation topic according to a newly published chemical research paper, Li et al. (2025). It is worth noticing that this paper was published in 2025, so it is impossible for ChemDFM-R to have encountered this paper during either the pretraining of the base model, Qwen2.5-14B, or our domain-pretraining process. To demonstrate the potential of ChemDFM-R, we assume the role of a researcher seeking a new research inspiration and engage a dialogue with the model about the subfield to which this paper belongs.
@@
First, as a researcher seeking inspiration, we first ask ChemDFM-R to introduce the oxygen evolution reaction (OER) briefly, including the definition and mechanism of it. Although the answer itself is a relatively rigid response to the question, the model’s rationale reveals interesting insights. In its rationale, ChemDFM-R mentions that OER requires a transition metal oxide catalyst, which often presents valuable opportunities for further research. Therefore, we further ask it to give more detailed recommendations for the catalysts. As a response, ChemDFM-R proposes multiple different oxides, which is a wide range for us to dig into. However, in its rationale, ChemDFM-R itself says that “iridium oxide (IrO2) is particularly notable for its …”, which is very inspiring. Since it is well known that iridium-based compounds are often very expensive, a natural follow-up question arises: how can we optimize this catalyst to reduce its cost while maintaining its catalytic performance? Surprisingly, ChemDFM-R manages to propose the initial ideas that align closely with the ideas presented in Li et al. (2025), “forming mixed metal oxides or heterostructures” and “optimizing the morphology and surface area of the catalysts”. At this point, a broad research direction has taken shape.
@@
It is worth noticing that nearly all the inspirations are drawn from the rationale of ChemDFM-R, which demonstrates the significance and value of ChemDFM-R’s ability to generate reasoning. This example shows that, with the enhanced chemical knowledge and strong chemical reasoning capabilities, ChemDFM-R has the potential to facilitate reliable human-AI collaboration, and so that advancing AI-driven research and applications.
@@ 5 Related Work
5.1 General Domain Reasoner LLMs
@@
Shortly after the emergence of LLMs, their remarkable emergent reasoning capabilities were discovered by Kojima et al. (2022) and explored by works such as ToT (Yao et al., 2023) and PAL (Gao et al., 2023). Recently, OpenAI-o1 (Jaech et al., 2024) followed by DeepSeek-R1 (Guo et al., 2025) and Kimi K1.5 (Team et al., 2025) demonstrated the prowess of reasoning models and the method to enhance LLMs’ reasoning capabilities using reinforcement learning-based pipelines. Subsequently, many studies have focused on improving the reasoning capabilities of models in various general domains, primarily in mathematics and coding. For example, Shao et al. (2024) and Zhang et al. (2024b) have further proven and discussed the effectiveness of reinforcement learning in terms of enhancing models’ reasoning capabilities, while Dou et al. (2024) and Zhang et al. (2025c) have explored better reward functions in mathematics and coding.
Despite the active research on reasoner models in general domains, there is little research focused on developing chemical reasoner LLMs. Considering that chemical reasoning requires domain-specific atomized knowledge and intrinsic logic, enhancing chemical reasoning capabilities demands specialized and targeted training. In this work, we designed a relatively comprehensive reward system for various chemical tasks to enable reinforcement learning tailored to chemical problems. Additionally, we developed an effective distillation data acquisition method to mitigate the knowledge gap of general LLMs, allowing us to leverage the power of general-domain reasoner LLMs to prevent the early unstable cold start phase of reinforcement learning.
5.2 Chemical LLMs
@@
The specialization of LLMs has become one of the most popular research areas after the emergence of LLMs in the general domain, including the development of chemical LLMs. LlaSMol (Yu et al., 2024) and Mol-Instruct (Fang et al., 2024) construct a chemical instruction tuning dataset and develop models that could excel on multiple chemical tasks, while extensive training with only chemical tasks has led to a substantive loss of natural language capabilities and task generalization ability in these models. Shortly after, Zhang et al. (2024a) leveraged high-quality instruction tuning and developed ChemLLM, which has acquired advanced chemical capabilities while retaining a considerable level of general language abilities. Furthermore, ChemDFM (Zhao et al., 2025b) achieved stronger chemical and generalization capabilities through domain pretraining and instruction tuning with both chemical data and general-domain data. It is worth noticing that all data used in previous work is from open-sourced databases or benchmarks. They usually directly model the correlation between molecules or reactions and the terminal phenomena, such as molecular properties or reaction yields, while the factors that truly have a direct impact are the functional groups within the molecules. Without this atomized chemical knowledge, it would be arduous for current chemical LLMs to generate rationales that align with chemical intuition and principles. Therefore, in this work, we tackle this issue by constructing a function-group-centric domain pretraining corpus to introduce this atomized chemical knowledge to LLMs. Recently, there has been some work starting to build reasoning models for specific chemical tasks, such as retrosynthesis (Zhang et al., 2025a), but a general chemical reasoner model is still absent.

### v3

@@ 4 Experimental Results
We first evaluate ChemDFM-R with multiple baselines on benchmarks specifically designed for assessing the chemical capabilities of LLMs (§˜4.1). Then, we validate the effectiveness of our methods by ablation studies (§˜4.2). Finally, to highlight the importance and effectiveness of chemical reasoning, we demonstrate its capability in facilitating reliable human-AI collaboration (§˜4.3).
@@ 4.1 Benchmark Evaluation
We evaluate ChemDFM-R and the baseline models on two of the most popular and comprehensive benchmarks specifically designed for assessing the chemical capabilities of LLMs: SciKnowEval (feng2024sciknoweval) and ChemEval (huang2024chemeval). Given the large number of tasks included in SciKnowEval (19 tasks) and ChemEval (36 tasks), to facilitate fair and clear comparison, we categorized the tasks into three groups: text-centric, molecule-centric, and reaction-centric tasks. Details of the task categorization are provided in the Appendix G.
@@ 4.1.1 Performance Comparison with Baselines
Table 2: Benchmark results on SciKnowEval and ChemEval. “mol.” stands for “molecule” and “react.” stands for “reaction”. Qwen-2.5-14B represent Qwen2.5-14B-Instruct. The numbers in the table represent the average performance over five inference runs, with the values in parentheses indicating the standard deviation. The best performance for each task is indicated using boldface. * We use RPS (peng2025surveyspeechlargelanguage) to balance the different scales of the scores on different tasks in the ChemEval benchmark.
@@
Qwen2.5-14B 77.0(0.11) 35.7(0.23) 72.1(0.24) 61.3(0.12) 79.6(1.43) 24.8(0.63) 46.8(1.81) 57.7(0.72)
ChemDFM-I 78.8(0.38) 50.6(0.40) 91.1(0.70) 69.7(0.23) 81.2(2.48) 68.4(3.56) 54.6(5.19) 70.7(1.10)
ChemDFM-R 76.7(0.18) 51.1(0.56) 93.8(0.37) 69.1(0.27) 78.3(2.38) 83.5(1.89) 58.5(2.61) 73.8(1.83)
First, we show the effectiveness of our training pipeline by comparing the performances of ChemDFM-R with those of 1) Qwen2.5-14B-Instruct (yang2024qwen25), which is the general-domain instruction tuning model of Qwen2.5-14B, and 2) ChemDFM-I, which incorporates atomized chemical knowledge enhancement but precedes the stage of chemical rationale learning. The quantitative results are illustrated in Table 2, while examples of the ChemDFM-R’s rationales are analyzed in Appendix F.
As showcased in Table 2, ChemDFM-R consistently outperforms Qwen2.5-14B-Instruct on both SciKnowEval and ChemEval, demonstrating that our specialization pipeline has successfully improved the model’s chemical capabilities. Specifically, the performances on text-centric tasks remain almost intact while those on molecule-centric and reaction-centric tasks increase significantly. This proves that our method manages to improve the chemical capabilities of LLM while largely maintaining its abilities in understanding natural language.
Moreover, we also evaluated ChemDFM-I to illustrate the contributions of the different stages in our training pipeline. Results show that the atomized chemical knowledge enhancement stage consistently improves performance across all task categories, while the chemical rationale learning stage further strengthens performance on molecule- and reaction-centric tasks. We attribute this phenomenon to two factors. First, the molecule- and reaction-centric tasks typically demand more domain-specific chemical reasoning, such as molecular property prediction or retrosynthesis analysis. In contrast, text-centric tasks rely more on natural language understanding, such as chemical named entity recognition and literature question-answering. As a result, learning chemical reasoning over molecules and reactions provides limited benefit to these text-focused tasks. Second, for the sake of answer verifiability, the reinforcement learning tasks do not include purely text-based tasks, which may adversely affect the model’s text reasoning ability. Incorporating text-related tasks into the RL stage through joint training might help preserve performance on text-based tasks.
Furthermore, Figure 4 illustrates the performance changes across individual tasks. The results clearly show that most tasks benefit from our training pipeline, especially the molecule-centric tasks and reaction-centric tasks. Moreover, the two training stages provide complementary gains across different tasks, enabling the final model to achieve superior results on a broader range of tasks. Notably, among the tasks where ChemDFM-R does not surpass Qwen2.5-14B-Instruct, a substantial proportion involves numerical prediction, such as Yield Extraction, Molecular Property Calculation, and Product Yield Prediction. In fact, almost all the molecule-centric and reaction-centric tasks where ChemDFM-R falls short of Qwen2.5-14B-Instruct are those involving numerical reasoning and prediction. This pattern suggests that the numerical calculation and prediction abilities of ChemDFM-R are relatively weak, highlighting a potential direction for further improvements.
@@ 4.1.2 Performance Comparison with Advanced LLMs
To further demonstrate the prowess of ChemDFM-R, we compare it with three sets of models: 1) existing chemical LLMs, including MolInst (fang2023mol), ChemLLM (zhang2024chemllm), ChemDFM (zhao2025chemdfm), and ether0 (narayanan2025training); 2) advanced non reasoning LLMs in the general domain, including GPT-4o (hurst2024gpt4o) and Qwen3-14B (no think) (yang2025qwen3); 3) advanced reasoning LLMs in the general domain, including DeepSeek-R1 (guo2025deepseek), Qwen3-14B (think) (yang2025qwen3), and o4-mini (openai_o3_o4mini). The experimental results are illustrated in Table 3. For detailed performances of individual tasks and more evaluations on FGBench (liu2025fgbench) and ChemCOTBench (li2025beyond), please refer to Appendix G.
Table 3: Benchmark results on SciKnowEval and ChemEval. “mol.” stands for “molecule” and “react.” stands for “reaction”. The best performance for each task is indicated using boldface, while the second-best is indicated using underline. * We use RPS (peng2025surveyspeechlargelanguage) to balance the different scales of the scores on different tasks in the ChemEval benchmark.
@@
MolInst 69.4 38.4 41.7 55.0 47.3 15.6 8.9 29.1
ChemLLM-20B-DPO 73.6 38.4 88.1 62.1 64.5 25.4 17.5 42.2
ChemDFM-13B-v1.0 70.2 42.6 85.0 61.6 61.5 42.3 34.0 49.2
ChemDFM-8B-v1.5 72.2 45.6 86.9 63.9 26.7 12.6 18.3 21.1
ether0 38.7 33.7 66.0 39.8 9.4 14.3 11.2 11.0
GPT-4o 76.1 44.6 44.6 61.2 82.7 35.0 52.9 63.3
Qwen3-14B (no think) 76.5 34.5 80.2 61.4 81.5 25.4 26.9 53.1
DeepSeek-R1 80.1 30.8 83.0 62.2 81.2 19.7 47.9 57.6
Qwen3-14B (think) 86.6 40.8 88.5 69.9 79.0 24.7 40.7 55.7
o4-mini 81.3 57.0 97.3 74.0 78.4 60.7 53.6 67.3
ChemDFM-R 76.7 51.1 93.8 69.1 78.3 83.5 58.5 73.8
As shown in Table 3, ChemDFM-R significantly outperforms both the general-domain LLMs and domain-specific LLMs of similar size, especially in the molecule-centric and reaction-centric tasks. Specifically, on ChemEval, it even considerably surpasses Qwen3-14B, the next-generation model in the same series as our base model Qwen2.5-14B. When compared to cutting-edge LLMs, ChemDFM-R achieves better performance than GPT-4o and DeepSeek-R1, while demonstrating competitive results relative to o4-mini on SciKnowEval. Considering the tiny size of our model, this result strongly demonstrates the prowess of ChemDFM-R and the effectiveness of our specialization process.
4.2 Ablation Study
We conducted two sets of ablation studies, which verify the two key components of ChemDFM-R’s training pipeline, respectively: 1) atomized chemical knowledge enhanced domain pretraining (demonstrated in Appendix H); 2) mixed-source distillation (demonstrated below).
To validate the effectiveness of our newly designed mixed-source distillation method, we conduct an ablation study by gradually simplifying the composition of the distillation dataset. The results are shown in Table 4. The results prove that the traditional distillation method (Row 2) struggles to achieve positive impacts on performance in the chemical domain. It even underperforms the “zero” method (Row 1) proposed by Deepseek-R1, where there is no distillation stage before reinforcement learning. With the help of data sampled from the instruction-tuning dataset to maintain chemical capabilities and knowledge (Row 3), the model’s performance gets boosted significantly. Moreover, the pseudo-reasoning data further help the model to achieve higher performance (Row 4), which corresponds to the final setting of our proposed mixed-source distillation method.
Table 4: Ablation study results on SciKnowEval and ChemEval. “No Thinking” denotes the data source of the instruction-tuning dataset, “Pseudo” denotes the pseudo-reasoning data source, and “Distilled” denotes the teacher’s rationale data source. The best performance for each task is indicated using boldface. * We use RPS (peng2025surveyspeechlargelanguage) to balance the different scales of the scores on different tasks in the ChemEval benchmark.
Data Source for Distillation SciKnowEval ChemEval*
No Thinking Pseudo Distilled text mol. react. all text mol. react. all
77.7 49.9 94.2 69.2 76.4 70.9 59.9 70.4
75.5 49.6 87.9 67.2 72.0 68.8 63.6 68.9
77.4 50.6 92.4 69.1 78.0 83.1 61.8 74.5
76.8 52.0 94.5 69.5 80.3 84.5 61.3 75.8
@@ 4.3 Reliable Human-AI Collaboration
Another important advantage of ChemDFM-R’s reasoning capability is that it allows humans to verify the correctness of answers, identify and correct errors, and discover new insights or perspectives. This enables more practical, reliable, and flexible human-AI collaboration. In this section, we first present human evaluation scores of the rationales generated by different reasoning models, which highlight the advantages of ChemDFM-R in facilitating more practical, reliable, and user-friendly human–AI interaction. We then demonstrate the prowess of ChemDFM-R in enabling reliable human–AI collaboration through illustrative examples. Specifically, an example is showcased in Figure 5, while more examples are illustrated in Appendix J.
4.3.1 Human Assessment of Rationale Quality
To assess the quality of rationales under practical situations, we constructed ten graduate-level questions based on recent publications from several influential chemistry journals. The questions cover different major subfields of chemistry, including organic chemistry, inorganic chemistry, materials chemistry, analytical chemistry, and polymer chemistry. Then, different reasoning LLMs are leveraged to solve these questions through human-AI interactions. Five graduate-level chemistry experts were hired to evaluate these interactions across five dimensions. The evaluation results are demonstrated in Table 5, while the questions and metrics are introduced in Appendix I. From these results, we draw three main conclusions:
First, our model outperforms all baselines, including DeepSeek-R1, in both chemical correctness and answer accuracy, indicating that it possesses a stronger grasp of chemical knowledge. In particular, our model shows a clearer advantage on SMILES-related tasks, demonstrating a more precise understanding of molecular structures.
Second, our model achieves significantly higher scores in effective information density. Unlike models such as Qwen3-14B and DeepSeek-R1, which often generate extremely lengthy reasoning chains exceeding 1,000 tokens, our model produces concise yet informative reasoning. This considerably reduces the burden on human users when verifying correctness or identifying errors.
Finally, our model performs slightly lower than Qwen3-14B and DeepSeek-R1 on analytical coverage. This is partially because their extremely lengthy reasoning chains allow them to enumerate many possible considerations. It is also true that our model commonly focuses on one or a few major possibilities rather than adequately enumerating all potential factors. This highlights a productive direction for future enhancement of our model’s analytical breadth.
Table 5: Human assessment of the quality of rationales from different reasoning LLMs. Please refer to Appendix I for the introduction of the five metrics and the detailed questions used in this assessment. As the rationales of o4-mini are not accessible, we only use the answer provided by o4-mini. All metrics are higher-is-better.
Models Chemical Answer Analytical Reasoning Effective
Correctness Accuracy Coverage Coherence Information Density
ether0 1.56 1.60 1.40 1.76 1.56
DeepSeek-R1 4.10 3.88 4.34 4.42 3.62
Qwen3-14B (think) 3.56 3.92 4.24 3.98 3.00
o4-mini 3.94 3.96 3.84 4.16 4.12
ChemDFM-R 4.52 4.28 3.94 4.62 4.70
@@ 4.3.1 Human Assessment of Rationale Quality
Figure 5: An example of reliable human-AI collaboration using ChemDFM-R. This conversation is inspired by li2025achieving. This example illustrates a process of research inspiration seeking with the help of ChemDFM-R.
4.3.2 Practical Cases of Human-AI Collaboration
@@
To make our example more realistic, we design our conversation topic according to a newly published chemical research paper, li2025achieving. It is worth noticing that this paper was published in 2025, so it is impossible for ChemDFM-R to encounter this paper during either the pretraining of the base model, Qwen2.5-14B, or our domain-pretraining process. To demonstrate the potential of ChemDFM-R, we assume the role of a researcher seeking a new research inspiration and engage the model in a dialogue about the subfield to which this paper belongs.
@@
First, as a researcher seeking inspiration, we ask ChemDFM-R to introduce the oxygen evolution reaction (OER) briefly, including the definition and mechanism of it. Although the answer itself is a relatively rigid response to the question, the model’s rationale reveals interesting insights. In its rationale, ChemDFM-R mentions that OER requires a transition metal oxide catalyst, which often presents valuable opportunities for further research. Therefore, we further ask it to give more detailed recommendations for the catalysts. As a response, ChemDFM-R proposes multiple different oxides, which is a wide range for us to dig into. However, in its rationale, ChemDFM-R itself says that “iridium oxide (IrO2) is particularly notable for its …”, which is very inspiring. Since it is well known that iridium-based compounds are often very expensive, a natural follow-up question arises: how can we optimize this catalyst to reduce its cost while maintaining its catalytic performance? Surprisingly, ChemDFM-R manages to propose the initial ideas that align closely with the ideas presented in li2025achieving, “forming mixed metal oxides or heterostructures” and “optimizing the morphology and surface area of the catalysts”. At this point, a broad research direction has taken shape.
@@
It is worth noticing that nearly all the inspirations are drawn from the rationale of ChemDFM-R, which demonstrates the significance and value of ChemDFM-R’s ability to generate reasoning. This example shows that, with the enhanced chemical knowledge and strong chemical reasoning capabilities, ChemDFM-R has the potential to facilitate reliable human-AI collaboration, thereby advancing AI-driven research and applications. More examples involving error correction and answer improvement with the help of rationales are demonstrated in Appendix J.
@@ 5 Related Work
General Domain Reasoning LLMs.
@@
Shortly after the emergence of LLMs, their remarkable reasoning capabilities were discovered by kojima2022large and explored by works such as ToT (yao2023tree) and PAL (gao2023pal). Recently, OpenAI-o1 (jaech2024openai) followed by DeepSeek-R1 (guo2025deepseek) and Kimi K1.5 (team2025kimi) demonstrated the prowess of reasoning models and the method to enhance LLMs’ reasoning capabilities using reinforcement learning-based pipelines. Subsequently, many studies have focused on improving the reasoning capabilities of models in various general domains, primarily in mathematics and coding. For example, shao2024deepseekmath and zhang2024o1 have further proven and discussed the effectiveness of reinforcement learning in terms of enhancing models’ reasoning capabilities, while dou2024stepcoder and zhang2025lessons have explored better reward functions in mathematics and coding.
Chemical LLMs.
@@
The specialization of LLMs has become one of the most popular research areas after the emergence of general-use LLMs, including the development of chemical LLMs. LlaSMol (yu2024llasmol) and Mol-Instruct (fang2023mol) construct a chemical instruction tuning dataset and develop models that could excel in multiple chemical tasks, while extensive training with only chemical tasks has led to a substantive loss of natural language capabilities and task generalization ability in these models. Shortly after, zhang2024chemllm leveraged high-quality instruction tuning and developed ChemLLM, which has acquired advanced chemical capabilities while retaining a considerable level of general language abilities. Furthermore, ChemDFM (zhao2025chemdfm) achieved stronger chemical and generalization capabilities through domain pretraining and instruction tuning with both chemical data and general-domain data. It is worth noticing that the data used in previous work overlooks the intrinsic chemical essence, which is crucial for LLMs to master reasoning with chemical intuition and principles. To tackle this issue, we construct a function-group-centric domain pretraining corpus to introduce such atomized chemical knowledge to LLMs. Recently, there has been pioneering work building chemical reasoning models for specific tasks, such as RetroDFM (zhang2025reasoningdrivenretrosynthesispredictionlarge) for retrosynthesis and multi-task specialist ether0 (narayanan2025training). Although their models exhibit strong reasoning capabilities and favorable performance on the training tasks, they often lack generalization ability. Consequently, they cannot accommodate the complex scenarios and demands that arise in real human–AI interactions, making it difficult for them to meet the requirements of a practical chemistry research assistant.